# A Survey on Employing Large Language Models for Text-to-SQL Tasks

### v1

@@ 1. Introduction
Text-to-SQL parsing is a well-established task in the field of natural language processing (NLP). Its purpose is to convert natural language queries into SQL queries, closing the gap between non-expert users and database access. To illustrate, imagine a table named cities with three columns: city_name (type: string), population (type: integer), and country (type: string). If we are given the natural language query ”Find all the cities with a population greater than 1 million in the United States,” the Text-to-SQL parsing technique should automatically generate the correct SQL query: ”SELECT city_name FROM cities WHERE population ¿ 1000000 AND country = ’United States’.” Researchers have made substantial progress in this area. Initially, template-based and rule-based methods(Zelle and Mooney, 1996; Li and Jagadish, 2014) were employed. These approaches involved creating SQL templates for various scenarios. While template-based methods showed promise, they required significant manual effort. With the rapid advancement of deep learning, Seq2Seq(Sutskever et al., 2014) methods have emerged as the mainstream approach. Seq2Seq(Sutskever et al., 2014) models provide an end-to-end solution, directly mapping natural language input to SQL output, which eliminates the need for intermediate steps like semantic parsing or rule-based systems. Among Seq2Seq(Sutskever et al., 2014) methods, pre-trained language models(PLMs), which serve as predecessors to large language models (LLMs), show promise in text-to-SQL tasks. Benefiting from the rich linguistic knowledge in large-scale corpora, PLMs become the state-of-the-art (SOTA) solution during that period(Qin et al., 2022).
@@
As model sizes and training data continue to grow, pre-trained language models (PLMs) naturally evolve into large language models (LLMs), exhibiting even greater power. Due to the scaling law and their emergent capabilities, LLMs have made substantial contributions across diverse domains, including chatbots(Dam et al., 2024), software engineering(Hou et al., 2023), and agents(Wang et al., 2024), etc. The remarkable capabilities of LLMs have prompted research into their application for text-to-SQL tasks. Figure  1 shows the evolution of mainstreaming approaches(Lin et al., 2020; Mellah et al., 2021; Rubin and Berant, 2020; Scholak et al., 2021; Qi et al., 2022; Li et al., 2023b; Pourreza and Rafiei, 2024; Dong et al., 2023b; Gao et al., 2023a; Gan et al., 2021c) in the Text-to-SQL domain, represented by the Execution Accuracy in SPIDER test dataset(Yu et al., 2018). Current literature mainly focuses on two main approaches of LLMs: prompt engineering and fine-tuning. Prompt engineering leverages methods like RAG, few-shot learning, and reasoning, while fine-tuning involves training an LLM on task-specific data. There exists a trade-off between prompt engineering and fine-tuning. Usually, prompt engineering demands less data but may lead to suboptimal results, while fine-tuning can enhance performance but necessitates a larger training dataset.
Figure 1. The evolution of Text-to-SQL approach over time
@@
This survey aims to provide a comprehensive overview of the use of LLMs in text-to-SQL tasks, including:
  • LLMs and text-to-SQL tasks: We provide a concise overview of LLMs and text-to-SQL tasks. Regarding LLMs, we delve into the factors contributing to their success, namely the scaling law(Kaplan et al., 2020) and emergent capability(Wei et al., 2022a). Additionally, we review the predominant approaches employed in LLM applications, namely prompt engineering and finetuning. As for text-to-SQL tasks, we introduce the Seq2Seq(Sutskever et al., 2014) method paradigm.
@@
2. Background
2.1. Large Language Models
Large Language Models(LLMs)(Chang et al., 2024) mainly refer to transformer-based(Vaswani et al., 2017) neural language models which contain billions of parameters and are pre-trained on large-scale corpus. LLMs like GPT-4(Achiam et al., 2023) and LLaMA(Touvron et al., 2023a) have shown an impressive effect on a wide range of natural language tasks such as translation, summary, intent detection, coding, etc. LLMs’ success is due to scaling law(Kaplan et al., 2020) and the emergent capability(Wei et al., 2022a). Scaling law(Kaplan et al., 2020) points out that language models’ performance on one specific downstream task has a power law relationship with three factors: the size of parameters, dataset size, and training time. The emergent capability(Wei et al., 2022a) indicates that some abilities may not be present in regular models but can arise in much larger models. One example of emergent capability is few-shot learning(Brown et al., 2020), meaning that large language models can complete the downstream tasks with several proper task demonstrations in the prompt without further training. Another example is the instruction following ability, which by finetuning on a mixture of tasks phrased as instructions, language models have been shown to respond appropriately to instructions describing an unseen task.
As of the time the authors write this paper, there are mainly two kinds of LLMs that are widely used. One type is the closed-source large language model, represented by the OpenAI GPT series(Liu et al., 2023b; Floridi and Chiriatti, 2020) and Google Gemini(Team et al., 2023), which generally can be accessed through official APIs. They typically have hundreds of billions or even trillions of parameters with strong emergence capabilities; Another type is the publicly available large language model, represented by LLaMA(Touvron et al., 2023a), which can be downloaded locally by users for private deployment or fine-tuning.
2.2. Prompt Engineering and Fine-Tuning
As for practice, there are two main streams of LLMs’ application to downstream tasks, prompt engineering and fine-tuning.
The representative approaches of prompt engineering are Retrieval Augmented Generation(RAG)(Gao et al., 2023b), few-shot learning(Brown et al., 2020), and reasoning. RAG(Gao et al., 2023b) means retrieving related and latest documents using the inputs and constructing the prompt using both concatenated documents and inputs. With the help of RAG, LLMs can acquire the latest knowledge, improve the interpretability of the response, and avoid hallucinations to some extent. As mentioned above, few-shot learning(Brown et al., 2020) is another impressive approach for LLM to adapt to new tasks quickly. Finally, the reasoning capability(Wei et al., 2022b; Zhou et al., 2022; Yao et al., 2024) of LLM has been discovered by researchers which enables LLM to ”think about” some problems sequentially and logically. The other application of LLM is fine-tuning, which means post-training LLM with an additional amount of specific domain corpus.
The second approach, fine-tuning, addresses the privacy challenge while raising the cost due to the requirement of high computing power to tune an LLM. There are several approaches to fine-tuning an LLM, including Supervised Fine-Tuning(SFT), Reinforcement Learning from Human Feedback(RLHF)(Bai et al., 2022), and Parameter-Efficient Fine-Tuning(PEFT)(Han et al., 2024). In SFT, the finetuning data is collected from a set of responses validated beforehand, which is the main difference to unsupervised techniques. Instruction tuning(Zhang et al., 2023b) is a typical method of SFT, which constructs a collection of instruction-formatted instances, and then directly fine-tunes a pretrained model in a supervised manner. RLHF combines the advantages of reinforcement learning(Wang et al., 2022b) and the pre-trained large model to involve human feedback. RLHF mainly includes the following three key steps: supervised fine-tuning, reward model training, and reinforcement learning fine-tuning. Since it would be costly to perform the full parameter tuning, PEFT chooses to reduce the number of trainable parameters while retaining good performance. Among the methods in PEFT, Low-Rank Adaptation(LoRA) of LLMs(Hu et al., 2021) imposes the low-rank constraint for approximating the update matrix at each dense layer; Prefix tuning(Li and Liang, 2021) prepares a series of prefixes, which are trainable continuous vectors and can be considered as virtual token embeddings, for each Transformer layer in the pre-trained large model; Prompt tuning(Lester et al., 2021) focuses on adjusting the prompts that guide the model’s response instead of modifying the deep structural weights of the model.
2.3. Text-to-SQL
Text-to-SQL is a classic natural language processing(NLP) problem about translating natural language into SQL statements(Katsogiannis-Meimarakis and Koutrika, 2023). Existing deep-learning methods focus on schema linking, jointly encoding natural language questions(NLQ) together with database schema and decoding to predict the target SQL.
On the schema-linking side, the current literature tries to link parts of NLQ to the elements in the database schema. IRNet(Jha et al., 2019) uses all n-grams of several lengths in the questions to find all possible query candidates. ISESL-SQL(Liu et al., 2022a) iteratively builds a semantic enhanced schema-linking graph between question tokens and database schemas.
On the encoder side, some work chooses to separate NLQ and column encoding(Zhong et al., 2017b; Xu et al., 2017). HydraNet(Lyu et al., 2020) encodes NLQ with each column separately. Others use graphs to represent tables, columns, and their relationship(Wang et al., 2019; Bogin et al., 2019).
On the decoder side, the simplest approach is generating the predicted SQL as a sequence of words(Zhong et al., 2017b). Ryansql(Choi et al., 2021) proposes a sketch-based slot-filling approach and adds grammar rules to instruct the output. As a great deep learning synthesizer in the text-to-sql domain, Resdsql(Li et al., 2023b) proposes a ranking-enhanced encoding and skeleton-aware decoding framework to decouple schema linking and skeleton parsing.
@@ 4.1. Basic Structure of Prompt Engineering in Text2SQL
Others In addition to the four commonly used representations mentioned above, the work in (Arora et al., [n. d.]) concatenated the prompt words of API Documents and SELECT X as its own prompt words; (Arora et al., 2023) added random sampling of categorical columns and range information for numerical columns after CREATE TABLE statements; Some works (Zhang et al., 2023a; Dong et al., 2023b; Nan et al., 2023; Sun et al., 2023; Xie et al., 2024) separated primary or foreign keys from database descriptions and described them separately, attempting to increase LLM’s attention to primary and foreign keys; The work of (Kothyari et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b; Guo et al., 2023b; Sun et al., 2023) uses the form of Table.Column to list database schema, which is usually used in conjunction with other supplementary knowledge filling methods about schema; (Chang and Fosler-Lussier, 2023) modified the form of SELECT X to list column-wise data, for the purpose of explicitly enumerate categorical value.
@@ 4.2.1. Schema Knowledge
Whether researchers are predicting research datasets in models, or data engineers are writing SQL in real-world scenarios to solve Text2SQL tasks, they all face the challenge of having too many databases, too long schema, and difficulty in identifying effective table and column information. The major approach to solving this problem is to focus the model’s attention on schema related to the problem and ignore irrelevant schema, known as the schema linking method. This method is considered to incorporate schema knowledge into the prompt words. Many works (Zhang et al., 2023a; Dong et al., 2023b; Kothyari et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Nan et al., 2023; Tai et al., 2023; Wang et al., 2023b; Sui et al., 2023; Li et al., 2024b) proposed their customized approach on schema linking. Research(Zhang et al., 2023a) represents all schema items in the form of Table.Column, and then calculates their similarity to the current problem separately. The tables and columns with the highest similarity are added to the prompt; At the same time, the similarity score between each table and the current problem is also calculated, and the table with the highest score is added to the prompt. These most relevant schema items will guide LLM in the Chain-of-Thought process of generating SQL when solving new problems. (Dong et al., 2023b) built three key modules on the Text2SQL task based on ChatGPT, and used a two-step schema linking in one of its modules. Specifically, the first step is to use ChatGPT to recall the tables most relevant to the issue, and the second step is to use ChatGPT to recall the columns most relevant to the question. CRUSH4SQL (Kothyari et al., 2023) first requires LLM to use only the original natural language problem as the prompt, add relevant examples, and generate a minimal database schema that can be used to solve the problem in a ”hallucinatory” manner. Then, using this ”hallucinatory” schema as an index, similarity retrieval is performed on the real schema to obtain subsets related to the problem in the real schema. DIN-SQL(Pourreza and Rafiei, 2024) utilizes the SPIDER dataset to generate examples as prompt words in the schema-linking scenario. The author organizes the examples in the form of Chain-of-Thought(Wei et al., 2022b), including natural language questions in SPIDER, keywords extracted from the questions, and table schema corresponding to the keywords. (Nan et al., 2023) has enhanced the semantics of the schema by representing the relationships between tables in the form of Entity Relationship, and specifically explaining which tables can be joined. (Tai et al., 2023) first breaks down Text2SQL into several sub-problems, and generates the corresponding most likely schema for each sub-problem. (Sui et al., 2023) adopts the Explain-Squeeze method, The Explain section uses LLM to answer schema-related questions step by step, including information dimensions in the query, identifying tables in the query, listing and explaining the columns in these tables, and locating the columns that appear in the query; The Squeeze section extracts the final answer from LLM’s answers to the above questions. (Li et al., 2024b) perform schema linking after generating the PreSQL statement in the first step, and then use the selected schema for prompt simplification, and finally use the simplified prompt to generate SQL for LLM.
@@ 4.3. Example Selection
If we want LLM pre-trained on general domain data to achieve good results in the Text2SQL domain, providing examples for the Text2SQL domain is a good solution, just like humans are good at imitating similar situations to make more accurate decisions. The LLM-related work in the field of Text2SQL is divided into two categories in the selection of examples: One kind of work (Liu et al., 2023a; Dong et al., 2023b; Hu et al., 2023; Wang et al., 2023a; Liu and Tan, 2023; Rajkumar et al., 2022; Chang and Fosler-Lussier, 2023; Sui et al., 2023) adopts zero-shot strategy, which does not include Text2SQL examples in prompt words, but focuses on other aspects like the table column structure of the database and dividing Text2SQL problems into multiple stages to solve; The other kind of work (Guo et al., 2023a; Arora et al., [n. d.]; Zhang et al., 2023a; Arora et al., 2023; Cheng et al., 2022; Li et al., 2024a; Kothyari et al., 2023; Pourreza and Rafiei, 2024; Nan et al., 2023; Tai et al., 2023; Gu et al., 2023; Wang et al., 2023b; Sun et al., 2023; Gao et al., 2023a; Xie et al., 2024; Li et al., 2024b) adopts few-shot strategy(Brown et al., 2020), which includes several Text2SQL examples in the design of prompt words. It is hoped that LLM can learn the correct pattern for generating SQL from the examples, which can be adopted in many subdivision stages of Text2SQL. Since zero-shot methods are usually only used to save the number of tokens consumed by LLM and reduce costs, the following will mainly introduce the existing work on the application of few-shot methods.
@@
Few-shot The few-shot method can be used to improve the effectiveness of nearly every corresponding step in the Text2SQL task. (Arora et al., [n. d.]; Li et al., 2024a; Sun et al., 2023) select the combination example of database schema, text query, and SQL statements which provides direct guidance for Text2SQL tasks. To reduce the number of tokens input into LLM and save costs, sometimes the database schema was removed in the examples (Nan et al., 2023; Gao et al., 2023a; Xie et al., 2024), while text query and SQL statements were retained. (Guo et al., 2023a, b) only select a few SQL statements and add prompt words to allow LLM to have a comprehensive understanding of SQL syntax, keywords, and possible solutions to the current problem. (Zhang et al., 2023a; Arora et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b) merge the text query, SQL statements, and inference processes into the example to guide LLM in inferring the current problem according to the given pattern. (Cheng et al., 2022) apply the few-shot method to the generation of its own BINDER-SQL, allowing LLM which is completely unfamiliar with the BINDER-SQL language to understand the common writing methods of the language through examples, laying the foundation for its subsequent API calls and SQL generation. Unlike providing examples related to SQL, (Kothyari et al., 2023) focuses on creating illusions for LLM when inferring schemas, with examples in prompt words including text queries and their corresponding nonreal schemas. Similar to using the few-shot method in schema prediction, (Pourreza and Rafiei, 2024) used partial text queries and corresponding real schema from the SPIDER dataset to form the examples in the prompt. (Gu et al., 2023) using prompt learning, divide SQL generation into two parts: SQL structure and SQL content, and use feed shot to solve them separately.

### v2

@@ 1. Introduction
Text-to-SQL parsing is a well-established task in the field of natural language processing (NLP). Its purpose is to convert natural language queries into SQL queries, bridging the gap between non-expert users and database access. To illustrate, imagine a table named cities with three columns: city_name (type: string), population (type: integer), and country (type: string). If we are given the natural language query ”Find all the cities with a population greater than 1 million in the United States,” the Text-to-SQL parsing technique should automatically generate the correct SQL query: ”SELECT city_name FROM cities WHERE population ¿ 1000000 AND country = ’United States’.” Researchers have made substantial progress in this area. Initially, template-based and rule-based methods(Zelle and Mooney, 1996; Li and Jagadish, 2014) were employed. These approaches involved creating SQL templates for various scenarios. While template-based methods showed promise, they required significant manual effort. With the rapid advancement of deep learning, Seq2Seq(Sutskever et al., 2014) methods have emerged as the mainstream approach. Seq2Seq(Sutskever et al., 2014) models provide an end-to-end solution, directly mapping natural language input to SQL output, which eliminates the need for intermediate steps like semantic parsing or rule-based systems. Among Seq2Seq(Sutskever et al., 2014) methods, pre-trained language models(PLMs), which serve as predecessors to large language models (LLMs), show promise in text-to-SQL tasks. Benefiting from the rich linguistic knowledge in large-scale corpora, PLMs become the state-of-the-art (SOTA) solution during that period(Qin et al., 2022).
@@
As model sizes and training data continue to grow, pre-trained language models (PLMs) naturally evolve into large language models (LLMs), exhibiting even greater power. Due to the scaling law and their emergent capabilities, LLMs have made substantial contributions across diverse domains, including chatbots(Dam et al., 2024), software engineering(Hou et al., 2023), and agents(Wang et al., 2024), etc. The remarkable capabilities of LLMs have prompted research into their application for text-to-SQL tasks. Figure  2 shows the evolution of mainstreaming approaches(Lin et al., 2020; Mellah et al., 2021; Rubin and Berant, 2020; Scholak et al., 2021; Qi et al., 2022; Li et al., 2023b; Pourreza and Rafiei, 2024; Dong et al., 2023b; Gao et al., 2023a; Gan et al., 2021c) in the Text-to-SQL domain, represented by the Execution Accuracy in SPIDER test dataset(Yu et al., 2018). Current literature mainly focuses on two main approaches of LLMs: prompt engineering and fine-tuning. Prompt engineering leverages methods like RAG, few-shot learning, and reasoning, while fine-tuning involves training an LLM on task-specific data. There exists a trade-off between prompt engineering and fine-tuning. Usually, prompt engineering demands less data but may lead to suboptimal results, while fine-tuning can enhance performance but necessitates a larger training dataset.
@@
This paper aims to provide a comprehensive survey of the use of LLMs in text-to-SQL tasks. We provide an overview of LLMs’ characteristics and advantages over traditional models, and introduce LLM-based Text2SQL in the following aspects:
@@
Figure 1. Framework of employing LLMs in Text-to-SQL
2. Overview
2.1. LLMs and LLM-based Text-to-SQL
Large language models(LLMs) have emerged as a milestone for natural language processing and machine learning. The concept of LLMs comes from the practice of continually scaling up the parameter size of pretrained language models (PLMs) and the volume of training data(Zhao et al., 2023), which results in surprising abilities, known as emergent abilities(Wei et al., 2022a), that are not found in smaller PLMs. One example of emergent capability is few-shot learning(Brown et al., 2020), meaning that LLMs can complete the downstream tasks with several proper task demonstrations in the prompt without further training. Another example is the instruction following ability(Wei et al., 2021), with which LLMs have been shown to respond appropriately to instructions describing an unseen task.
Due to the emergent abilities(Wei et al., 2022a) of LLMs and the basic operating principle of LLMs which gradually produce the next word that has the highest probability based on the input prompt(Zhao et al., 2023), prompt engineering becomes one of the two primary streams to apply LLMs to downstream tasks. The representative approaches of prompt engineering are Retrieval Augmented Generation(RAG)(Gao et al., 2023b), few-shot learning(Brown et al., 2020), and reasoning(Wei et al., 2022b; Zhou et al., 2022; Yao et al., 2024). The other stream of LLMs’ application to downstream tasks is called fine-tuning, which follows the ”pre-training and fine-tuning” learning paradigm of PLMs(Zhao et al., 2023) and aims to boost performance in specific domains and address privacy concerns. The general fine-tuning process mainly involves data preparation, pre-training model selection, model fine-tuning, and model evaluation.
Text-to-SQL, which is a challenging task in both the natural language processing(NLP) and database communities and involves mapping natural language questions on the given relational database into SQL queries, has also been revolutionized by the emergence of LLMs. We summarize the general framework of LLM-based Text-to-SQL systems in fig.1. Based on the two main streams of LLM applications, we categorize the methods used in LLM-based Text-to-SQL into two categories, namely prompt engineering and fine-tuning. As for prompt engineering methods, we typically design a well-structured prompt encompassing various components, such as task descriptions, table schemas, questions and additional knowledge, and utilize in-context learning and reasoning methods at the same time. As for fine-tuning methods, we typically generate or collect Text-to-SQL datasets, select appropriate pretrained LLMs and finetuning approaches such as LORA(Hu et al., 2021), and compare the test results before and after to understand the changes in the model’s performance.
2.2. Difference between traditional and LLM-based Text-to-SQL Approaches
Prior to the widespread of LLMs, there were two primary streams for text-to-SQL methods. One stream involves utilizing the sequence-to-sequence (Seq2Seq) model(Sutskever et al., 2014), in which an encoder is designed to capture the semantics of the natural language(NL) question and the corresponding table schema, and a decoder is employed to generate the SQL query token by token based on the encoded question representation. Some notable methods in this approach include IRNet(Jha et al., 2019), SQLNet(Xu et al., 2017), Seq2SQL(Zhong et al., 2017b), HydraNet(Lyu et al., 2020), Ryansql(Choi et al., 2021), Resdsql(Li et al., 2023b) and ISESL-SQL(Liu et al., 2022a). The other stream involves finetuning PLMs such as BERT(Devlin, 2018), which leverage the extensive knowledge present in large-scale text collections and have proved to be effective in enhancing the performance of downstream text-to-SQL parsing tasks.
We argue that there are two distinct aspects that differentiate traditional text-to-SQL approaches from LLM-based ones:
  • Novel Paradigm: Traditional text-to-SQL approaches necessitate training, whereas LLMs can often circumvent this requirement. Leveraging the instruction following ability(Wei et al., 2021) of LLMs, LLMs can accomplish the text-to-SQL tasks with the aid of appropriate instructions and information.
  • Uniform Architecture: According to a previous survey(Qin et al., 2022), the encoder and decoder in traditional approaches can be designed using diverse architectures such as LSTM(Hochreiter and Schmidhuber, 1997), Transformer(Vaswani et al., 2017), and even GNNs(Bogin et al., 2019). In contrast, LLMs adhere to a uniform transformer-based architecture, which not only enables easier scaling up but also facilitates a more streamlined implementation.
2.3. Why LLM-based Text-to-SQL
There has been a sharp increase in the number of LLMs employed for text-to-SQL tasks in recent times. After conducting a comprehensive survey of recent papers, we identify several key reasons for this trend, which are summarized below:
  • Enhanced Performance: Figure  2 illustrates the progression of mainstreaming approaches within the Text-to-SQL domain, as indicated by the Execution Accuracy in the SPIDER test dataset(Yu et al., 2018). As we can see, LLM-based methods have significantly improved the SOTA performance, demonstrating the remarkable capabilities of LLM-based methods. The listed references(Lin et al., 2020; Mellah et al., 2021; Rubin and Berant, 2020; Scholak et al., 2021; Qi et al., 2022; Li et al., 2023b; Pourreza and Rafiei, 2024; Dong et al., 2023b; Gao et al., 2023a; Gan et al., 2021c) support this observation.
  • Generalization Ability and Adaptability: As we mentioned above, LLMs have introduced a new paradigm, namely prompt engineering, which benefits from their ability to follow instructions(Wei et al., 2021), making LLMs easily transferable to different settings without additional training. Furthermore, the in-context learning ability(Zhao et al., 2023) of LLMs further enhances their generalization and adaptability capabilities, as they can learn from examples provided to seamlessly fit into various settings.
  • Future Improvements: LLM-based methods hold a promising future for advancements. Since the global community is prioritizing the enhancement of LLMs, efforts and resources are concentrated towards supporting research on LLMs. This includes scaling up LLMs, creating new prompting methods, generating high-quality and extensive datasets, and fine-tuning LLMs across various tasks. The progress made by the LLM community will undoubtedly and continuously drive LLM-based text-to-SQL methods to new SOTA.
Figure 2. The evolution of Text-to-SQL approach over time
@@ 4.1. Basic Structure of Prompt Engineering in Text2SQL
Others In addition to the four commonly used representations mentioned above, the work in (Arora et al., [n. d.]) concatenated the prompt words of API Documents and SELECT X as its own prompt words; (Arora et al., 2023) added random sampling of categorical columns and range information for numerical columns after CREATE TABLE statements; Some works (Zhang et al., 2023; Dong et al., 2023b; Nan et al., 2023; Sun et al., 2023; Xie et al., 2024) separated primary or foreign keys from database descriptions and described them separately, attempting to increase LLM’s attention to primary and foreign keys; The work of (Kothyari et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b; Guo et al., 2023b; Sun et al., 2023) uses the form of Table.Column to list database schema, which is usually used in conjunction with other supplementary knowledge filling methods about schema; (Chang and Fosler-Lussier, 2023) modified the form of SELECT X to list column-wise data, for the purpose of explicitly enumerate categorical value.
@@ 4.2.1. Schema Knowledge
Whether researchers are predicting research datasets in models, or data engineers are writing SQL in real-world scenarios to solve Text2SQL tasks, they all face the challenge of having too many databases, too long schema, and difficulty in identifying effective table and column information. The major approach to solving this problem is to focus the model’s attention on schema related to the problem and ignore irrelevant schema, known as the schema linking method. This method is considered to incorporate schema knowledge into the prompt words. Many works (Zhang et al., 2023; Dong et al., 2023b; Kothyari et al., 2023; Wang et al., 2023a; Pourreza and Rafiei, 2024; Liu and Tan, 2023; Nan et al., 2023; Tai et al., 2023; Wang et al., 2023b; Sui et al., 2023; Li et al., 2024b) proposed their customized approach on schema linking. Research(Zhang et al., 2023) represents all schema items in the form of Table.Column, and then calculates their similarity to the current problem separately. The tables and columns with the highest similarity are added to the prompt; At the same time, the similarity score between each table and the current problem is also calculated, and the table with the highest score is added to the prompt. These most relevant schema items will guide LLM in the Chain-of-Thought process of generating SQL when solving new problems. (Dong et al., 2023b) built three key modules on the Text2SQL task based on ChatGPT, and used a two-step schema linking in one of its modules. Specifically, the first step is to use ChatGPT to recall the tables most relevant to the issue, and the second step is to use ChatGPT to recall the columns most relevant to the question. CRUSH4SQL (Kothyari et al., 2023) first requires LLM to use only the original natural language problem as the prompt, add relevant examples, and generate a minimal database schema that can be used to solve the problem in a ”hallucinatory” manner. Then, using this ”hallucinatory” schema as an index, similarity retrieval is performed on the real schema to obtain subsets related to the problem in the real schema. DIN-SQL(Pourreza and Rafiei, 2024) utilizes the SPIDER dataset to generate examples as prompt words in the schema-linking scenario. The author organizes the examples in the form of Chain-of-Thought(Wei et al., 2022b), including natural language questions in SPIDER, keywords extracted from the questions, and table schema corresponding to the keywords. (Nan et al., 2023) has enhanced the semantics of the schema by representing the relationships between tables in the form of Entity Relationship, and specifically explaining which tables can be joined. (Tai et al., 2023) first breaks down Text2SQL into several sub-problems, and generates the corresponding most likely schema for each sub-problem. (Sui et al., 2023) adopts the Explain-Squeeze method, The Explain section uses LLM to answer schema-related questions step by step, including information dimensions in the query, identifying tables in the query, listing and explaining the columns in these tables, and locating the columns that appear in the query; The Squeeze section extracts the final answer from LLM’s answers to the above questions. (Li et al., 2024b) perform schema linking after generating the PreSQL statement in the first step, and then use the selected schema for prompt simplification, and finally use the simplified prompt to generate SQL for LLM.
@@ 4.3. Example Selection
If we want LLM pre-trained on general domain data to achieve good results in the Text2SQL domain, providing examples for the Text2SQL domain is a good solution, just like humans are good at imitating similar situations to make more accurate decisions. The LLM-related work in the field of Text2SQL is divided into two categories in the selection of examples: One kind of work (Liu et al., 2023a; Dong et al., 2023b; Hu et al., 2023; Wang et al., 2023a; Liu and Tan, 2023; Rajkumar et al., 2022; Chang and Fosler-Lussier, 2023; Sui et al., 2023) adopts zero-shot strategy, which does not include Text2SQL examples in prompt words, but focuses on other aspects like the table column structure of the database and dividing Text2SQL problems into multiple stages to solve; The other kind of work (Guo et al., 2023a; Arora et al., [n. d.]; Zhang et al., 2023; Arora et al., 2023; Cheng et al., 2022; Li et al., 2024a; Kothyari et al., 2023; Pourreza and Rafiei, 2024; Nan et al., 2023; Tai et al., 2023; Gu et al., 2023; Wang et al., 2023b; Sun et al., 2023; Gao et al., 2023a; Xie et al., 2024; Li et al., 2024b) adopts few-shot strategy(Brown et al., 2020), which includes several Text2SQL examples in the design of prompt words. It is hoped that LLM can learn the correct pattern for generating SQL from the examples, which can be adopted in many subdivision stages of Text2SQL. Since zero-shot methods are usually only used to save the number of tokens consumed by LLM and reduce costs, the following will mainly introduce the existing work on the application of few-shot methods.
@@
Few-shot The few-shot method can be used to improve the effectiveness of nearly every corresponding step in the Text2SQL task. (Arora et al., [n. d.]; Li et al., 2024a; Sun et al., 2023) select the combination example of database schema, text query, and SQL statements which provides direct guidance for Text2SQL tasks. To reduce the number of tokens input into LLM and save costs, sometimes the database schema was removed in the examples (Nan et al., 2023; Gao et al., 2023a; Xie et al., 2024), while text query and SQL statements were retained. (Guo et al., 2023a, b) only select a few SQL statements and add prompt words to allow LLM to have a comprehensive understanding of SQL syntax, keywords, and possible solutions to the current problem. (Zhang et al., 2023; Arora et al., 2023; Pourreza and Rafiei, 2024; Wang et al., 2023b) merge the text query, SQL statements, and inference processes into the example to guide LLM in inferring the current problem according to the given pattern. (Cheng et al., 2022) apply the few-shot method to the generation of its own BINDER-SQL, allowing LLM which is completely unfamiliar with the BINDER-SQL language to understand the common writing methods of the language through examples, laying the foundation for its subsequent API calls and SQL generation. Unlike providing examples related to SQL, (Kothyari et al., 2023) focuses on creating illusions for LLM when inferring schemas, with examples in prompt words including text queries and their corresponding nonreal schemas. Similar to using the few-shot method in schema prediction, (Pourreza and Rafiei, 2024) used partial text queries and corresponding real schema from the SPIDER dataset to form the examples in the prompt. (Gu et al., 2023) using prompt learning, divide SQL generation into two parts: SQL structure and SQL content, and use feed shot to solve them separately.